A particle is thrown vertically upwards with velocity u, find its velocity 1 second before it reaches the maximum height

Grade:11

1 Answers

Prasanna Vibhandik
6 years ago
as we have time taken to reach max height =u/a
we want to find velocity at (t-1)seconds
so by using first kinematical equation
v=u+a(t-1)=u+a(v/a-1)